# VISER: Visually-Informed System for Enhanced Robustness in Open-Set Iris Presentation Attack Detection

### 18 Mar 2026 v1

@@ Abstract
Human perceptual priors have shown promise in saliency-guided deep learning training, particularly in the domain of iris presentation attack detection (PAD). Common saliency approaches include hand annotations obtained via mouse clicks and eye gaze heatmaps derived from eye tracking data. However, the most effective form of human saliency for open-set iris PAD remains underexplored. In this paper, we conduct a series of experiments comparing hand annotations, eye tracking heatmaps, segmentation masks, and DINOv2 embeddings to a state-of-the-art deep learning-based baseline on the task of open-set iris PAD. Results for open-set PAD in a leave-one-attack-type out paradigm indicate that denoised eye tracking heatmaps show the best generalization improvement over cross entropy in terms of Area Under the ROC curve (AUROC) and Attack Presentation Classification Error Rate (APCER) at Bona Fide Presentation Classification Error Rate (BPCER) of 1%. Along with this paper, we offer trained models, code, and saliency maps for reproducibility and to facilitate follow-up research efforts.
Figure 1: Experimental pipeline and contributions. A deep learning iris presentation attack detection model trained without saliency guidance serves as the baseline. We evaluate multiple saliency modalities through saliency-guided training, and further assess foundation model-sourced embeddings using logistic regression, linear-SVM, and RBF-SVM classifiers. Performance is reported in terms of Area Under the ROC curve (AUROC) and Attack Presentation Classification Rate (APCER) at 1% Bonafide Presentation Classification Rate (BPCER), expressed as relative improvement over the baseline trained traditionally.
@@ 1 Introduction
The use of human saliency in saliency-guided deep learning has seen increased interest in domains like synthetic face detection, x-ray abnormality detection, and iris PAD. The use of human saliency is used to improve the quality of model’s feature maps [12], and to directly guide the learning process of the model [4]. These techniques have shown to increase a model’s ability to generalize to unseen data types leading to more trustworthy and accurate performance. However collecting human saliency can be an expensive and time-consuming. While attempts to maximize the amount of saliency samples needed for the effective training of models has been explored [9, 10], there is still the question of what form of human saliency is optimal for the best transfer of expert’s knowledge into deep learning.
This paper focuses on the latter point as we present VISER, a Visually Informed System for Enhanced Robustness in open-set iris PAD. VISER attempts to address the trade-off between the continuous but noisy signal that eye tracking provides and the information loss that occurs in translation from visual to motor saliency. Results show that eye tracking heatmaps improve over the D-NetPAD cross entropy baseline for average AUROC and APCER at BPCER of 1% with denoised eye tracking showing the best performance of the variants. The main contributions of this paper are as follows:
@@
  • RQ3:
    Does de-noising eye tracking gaze maps improve over standard eye tracking gaze maps in saliency-guided training?
@@ 3 Experimental Design
To address our research questions, we perform several experiments testing different saliency methods under the same conditions. To test open-set robustness, we train each configuration in a leave-one-attack-type out manner and perform 12 independent runs per attack type with a total of 7 attack types tested. We first conduct a baseline measurement of iris PAD classification on the DenseNet121-based D-NetPAD architecture [23, 48] using standard cross-entropy loss, from here this configuration is simply referred to as ”XENT”.
@@
Finally, with the emergence of capable foundation models in biometric tasks, we test three configurations using DINOv2-Base embeddings [41]. The details of these remaining configurations are outlined in the following sections.
@@ 3.1 Metrics
In accordance with ISO/IEC 30107 we report the Attack Presentation Classification Error Rate (APCER) at Bona Fide Presentation Classification Error Rate (BPCER) threshold of 1% for each configuration respectively [26], we also report the average Area Under the Receiver Operator Characteristic (AUROC) curve to provide easy comparison to other Iris PAD works. As we are comparing to the XENT baseline in all cases, specifically for a given model training method we report the difference between it’s performance and the XENT baseline performance. This does not change that a lower value is better for the APCER metric and a higher value is better the AUROC metric.
Our final ranking values are averages over all attack types compiled from multiple random variables: the variable for multiple runs per attack type left out, and the average over each category. Due to this we simply report point value ranking and exclude standard deviations. We define a better performance as lower (in the case of APCER) or higher (in the case of AUROC) ranking.
@@ 3.2 Dataset Sources
The images in our dataset reflect the following attack types and anomalies along with the bona fide samples and their sources: real irises [6, 34, 18, 55, 46, 33, 63, 60, 62], irises with textured contact lenses [34, 33, 63, 62], artificial (plastic eyes and glass prosthesis) eyes [34], diseased/unhealthy eyes [55, 54], printouts [34, 18, 32], synthetically-generated iris images [61], printouts of irises with textured contact lenses [34], and post-mortem iris images [58].
The eye tracking heatmaps were sourced from an eye tracking study conducted by Dowling et al. [15] who captured eye gaze heatmaps from a mixture of expert and non-expert examiners. Acquiring this variety of saliency allows us to perform the first study we are aware of comparing the performance of different human saliency types for saliency guided training in the task of iris PAD.
3.3 Saliency Methods
3.3.1 Segmentation Masks
3.3.2 Hand Annotations
3.3.3 Eye Tracking
Raw eye-tracking data collected from both expert and non-expert participants during iris image evaluations was processed into individual saliency heatmaps. For each participant normalized fixation coordinates were first corrected using per-participant polynomial remapping coefficients to account for calibration offset, and a Gaussian blur was applied to produce a continuous heatmap representation of visual attention. To produce a unified ground truth saliency map for each iris image, individual participant heatmaps were aggregated via a pixel-wise mean across all evaluators, expert and non-expert.
3.4 No Saliency: Foundation Model Baseline
Research has shown that foundation models show remarkable ability to distinguish both full-size and cropped iris samples in closed-set presentation attack detection settings [49, 50]. Specifically, Redwan et al. used a variety of foundation models to obtain feature embeddings for iris PAD images and trained three standard classifiers, logistic regression (LogReg) [8], support vector machine (SVM)[7], and support vector machine with radial basis function kernel (SVM-RBF) [47] on the extracted features. In the task of iris PAD for full size irises, [41] suggests that the clear winner was DINOv2-Base. In accordance with these findings we test open-set iris PAD using the same methodology laid out by Redwan et al. resulting in our final three methods, DINOv2+LogReg, DINOv2+SVM-Linear, and DINOv2+SVM-RBF.
@@ 3.4 No Saliency: Foundation Model Baseline
Table 1: Compact AUROC Results ($\uparrow$). We report the absolute AUROC for the DenseNet (XENT) baseline, while all other rows show the change ($\Delta$) relative to this baseline. Best results per column are bolded. Each column presents the results obtained for detecting this particular attack type, which was left out during training to guarantee an open-set iris PAD evaluation.
Method Printout Diseased Post Mortem Synthetic Contacts + Print Textured Contact Artificial Avg. $\Delta*$
Baseline Performance
DenseNet (XENT) 0.9169 0.6537 0.8012 0.6541 0.9717 0.7450 0.6550 0.7711
Segmentation Masks
Baseline Saliency-based +0.0087 -0.0052 +0.0385 -0.0511 -0.0212 +0.0345 +0.0232 +0.0039
Hand Annotations
High Entropy -0.0184 +0.0520 +0.0073 -0.0712 -0.0123 -0.0050 +0.0175 -0.0043
Equal Entropy -0.0010 +0.0847 -0.0098 -0.0122 -0.0021 +0.0274 +0.0277 +0.0164
Low Entropy -0.0001 -0.0077 +0.0294 -0.0160 -0.0384 +0.0140 -0.0157 -0.0049
Eye Tracking Saliency
Full Eye Tracking +0.0473 +0.0282 -0.0468 +0.0941 -0.0155 +0.0474 +0.2387 +0.0562
Initial Eye Tracking +0.0513 +0.0192 -0.0155 +0.0236 -0.0481 +0.0650 +0.2672 +0.0518
De-noised Full ET +0.0604 +0.0345 -0.0820 +0.0621 -0.0782 +0.0597 +0.2438 +0.0429
De-noised Initial ET +0.0627 +0.0574 -0.0645 +0.1090 -0.0453 +0.0372 +0.2692 +0.0608
Foundation Model (No Saliency)
DINOv2 + LogReg -0.0326 -0.1079 +0.0378 -0.0340 +0.0086 -0.0327 +0.1520 -0.0013
DINOv2 + SVM-Linear -0.0523 -0.1247 -0.0033 -0.0613 +0.0017 -0.0351 +0.1737 -0.0145
DINOv2 + SVM-RBF -0.0104 +0.0023 +0.1796 +0.1121 -0.0159 -0.0448 +0.1167 +0.0485
* Average is calculated across all attack types. Baseline row shows absolute AUROC.
Table 2: Same as in Table 1, except that compact $\Delta$ APCER @ BPCER=1% vs XENT Baseline ($\downarrow$) results are presented.
@@
High Entropy +0.0257 +0.0088 +0.0166 +0.0169 +0.0623 -0.0294 +0.0000 +0.0144
Equal Entropy +0.0879 -0.0535 +0.0423 +0.0094 +0.0173 -0.0341 +0.0088 +0.0112
Low Entropy -0.0403 +0.0079 +0.0282 +0.0094 +0.1676 -0.0082 +0.0103 +0.0250
@@
Full Eye Tracking -0.2765 -0.0210 +0.0153 -0.0664 +0.0824 -0.1198 -0.2353 -0.0888
Initial Eye Tracking -0.2701 -0.0473 +0.0743 -0.0430 +0.1753 -0.0834 -0.3815 -0.0822
De-noised Full ET -0.2893 -0.0412 +0.0717 -0.0927 +0.2194 -0.1127 -0.3552 -0.0857
De-noised Initial ET -0.3617 -0.0412 +0.0551 -0.1413 +0.1600 -0.0599 -0.3552 -0.1063
@@
DINOv2 + LogReg +0.1026 +0.0255 -0.1449 -0.0187 +0.0297 -0.0200 -0.0350 -0.0087
DINOv2 + SVM-Linear +0.1914 +0.0219 -0.0924 -0.0037 +0.1408 +0.0035 -0.0935 +0.0240
DINOv2 + SVM-RBF +0.2885 +0.0588 -0.5180 +0.0047 +0.2835 -0.0118 +0.0132 +0.0170
@@ 4 Results
We trained 12 independent models per attack-type-left-out for a total of 84 models for each method tested. For each run of 12 models, we report the average AUROC and APCER at BPCER 1% threshold. Table 1 shows these results expressed as each method’s improvement above the D-NetPAD XENT baseline where each row refers to a method and each column refers to the attack type that was left out during training with a final delta column summarizing the overall performance. Our findings indicate that initial phase eye tracking maps de-noised with HDBSCAN show the largest AUROC improvement over the XENT baseline at 0.061. Entire and initial phase eye tracking maps also showed improved AUROC performance over XENT with improvements of 0.056 and 0.052 respectively. The remaining saliency-based methods either failed to surpass the XENT baseline entirely or improved it by only 0.01. Thus the answer to RQ1 is affirmative, eye tracking saliency, particularly De-noised Initial Eye Tracking saliency, surpasses other saliency-based methods in open-set iris PAD.
@@
When comparing the results of foundation models against XENT, the results are mixed. While DINOv2 + SVM-RBF showed improvement over XENT with an average AUROC of 0.048, both DINOv2 + LogReg and DINOv2 + SVM-Linear failed to improve the baseline. Similar to table 1, table 2 shows the same methods and their improvements above the D-NetPAD XENT baseline of APCER at BPCER at 1% for each attack type left out and a final delta column summarizing the overall performance. When comparing the results of foundation models in this table, the results are poor. DINOv2 + LogReg was near even with XENT with an improvement of 0.0087 while DINOv2 + SVM-Linear and DINOv2 + SVM-RBF failed to improve the baseline. For the saliency-based methods, eye tracking methods showed better AUROC performance than segmentation masks and hand annotations as outlined in the previous section. When comparing the APCER at BPCER of 1%, Segmentation Masks and Hand Annotations fail to improve over the baseline. Eye tracking methods however show strong improvement over XENT with De-noised Initial ET improving by 0.1063 and Full Eye Tracking, De-noised Full ET, and Initial Eye Tracking improving by 0.0888, 0.0857, and 0.0852 respectively. With eye tracking methods improving XENT APCER by 0.08 or greater and only DINOv2 + LogReg improving the baseline at all, the answer to RQ2 is clear, saliency-guided training using eye tracking heatmaps outperform foundation models in open-set iris PAD.
4.3 Answering RQ3: Does de-noising eye tracking gaze maps improve over standard eye tracking gaze maps in saliency-guided training?
The results of RQ1 and RQ2 indicate that eye tracking improved the XENT baseline in AUROC and APCER at BPCER 1%, but how do the de-noised methods compare against the standard eye tracking maps? Overall, De-noised Initial ET was the best performing method above XENT baseline in both AUROC and APCER at BPCER 1%. For AUROC, De-noised Inital ET reported an improvement of 0.0608 compared to Initial Eye Tracking at 0.0518 while in APCER at BPCER 1%, De-noised Initial ET showed an improvement of 0.1063 compared to Initial Eye Tracking’s 0.0822. While the results show a clear advantage for de-noising the Initial Eye Tracking saliency maps, this is not the case for Full Eye Tracking maps. In both metrics, Full Eye Tracking outperforms De-noised Full ET with AUROC of $0.0562>0.0429$ and APCER at BPCER 1% of $0.0888>0.0857$ respectively. Therefore the answer to RQ3 is mixed, de-noising eye tracking gaze improves performance for Initial Eye Tracking saliency maps but is less effective on Full Eye Tracking saliency maps.

### 07 May 2026 v2

@@ Abstract
Human perceptual priors have shown promise in saliency-guided deep learning training, particularly in the domain of iris presentation attack detection (PAD). Common saliency approaches include hand annotations obtained via mouse clicks and eye gaze heatmaps derived from eye tracking data. However, the most effective form of human saliency for open-set iris PAD remains under-explored. In this paper, we conduct a series of experiments comparing hand annotations, eye tracking heatmaps, segmentation masks, and foundation model embeddings to a state-of-the-art deep learning-based baseline on the task of open-set iris PAD. Results for open-set PAD in a leave-one-attack-type out paradigm indicate that denoised eye tracking heatmaps show the best generalization improvement over cross entropy in Attack Presentation Classification Error Rate (APCER) at Bona Fide Presentation Classification Error Rate (BPCER) of 1%. Along with this paper, we offer trained models, code, and saliency maps for reproducibility and to facilitate follow-up research efforts.
Figure 1: Experimental pipeline and paper contributions. We start with a DenseNet121 architecture with no saliency as our baseline. Different forms of saliency are then tested on the architecture using saliency-guided training. Next foundation model embeddings are tested in SVM-Linear, SVM-RBF, and CNN classifiers. Finally, A/BPCER 1% results are expressed as +/- improvement over the DenseNet121 baseline.
@@ 1 Introduction
The use of human saliency in saliency-guided deep learning has seen increased interest in domains like synthetic face detection, x-ray abnormality detection, and iris PAD. Human saliency is used to improve the quality of model’s feature maps [12], and to directly guide the learning process of the model [4]. These techniques have shown to increase a model’s ability to generalize to unseen data types leading to more trustworthy and accurate performance. However, collecting human saliency can be expensive and time-consuming. While attempts to maximize the number of saliency samples needed for the effective training of models has been explored [9, 10], there is still the question of what form of human saliency is optimal for the best transfer of expert’s knowledge into deep learning. Finally, with the rapid evolution of highly-capable foundation models, it raises the question of whether experts and human saliency are needed to push the generalization abilities of PAD models.
Thus in this paper we present VISER, a Visually Informed System for Enhanced Robustness in open-set iris PAD. VISER attempts to address the trade-off between the continuous but noisy signal that eye tracking provides and the information loss that occurs in translation from visual to motor saliency. Results show that eye tracking heatmaps improve over the DenseNet121 cross entropy baseline for average APCER at BPCER of 1% with denoised eye tracking showing the best performance of the variants. The main contributions of this paper are as follows:
@@ 3 Experimental Design
To address our research questions, we perform several experiments testing different saliency methods under the same conditions. In selection of our baseline architecture, we looked to the latest Iris Liveness Detection Competition results [38]. While the winning algorithm is a proprietary model, the methods of the runner up team, D-NetPAD and SPAD [51, 43], are DenseNet121-based models and thereby motivated our choice of this backbone for our baseline. To further solidify this choice, we trained a closed-set DenseNet121 model on 20,000 full-size iris samples and a test set of 2,000 images for 12 independent runs achieving an average AUROC of 0.999 and an average AP/BPCER 1% of 0.0068. Motivated by this result, we set out to test open-set robustness. We train each configuration in a leave-one-attack-type out manner and perform 12 independent runs per attack type with a total of 7 attack types tested. We first conduct a baseline measurement of full-resolution iris PAD classification on the DenseNet121 architecture [23, 51] using standard cross-entropy loss, hereafter simply referred to as ”XENT”.
@@
Finally, with the emergence of capable foundation models in biometric tasks, we test three classifier configurations using embeddings from DINOv2-Large, DINOv3, and SigLIP2 [42, 52, 64]. The details of these remaining configurations are outlined in the following sections.
@@ 3.1 Metrics
In accordance with ISO/IEC 30107 we report the Attack Presentation Classification Error Rate (APCER) at Bona Fide Presentation Classification Error Rate (BPCER) threshold of 1% for each configuration respectively [26]. As we are comparing to the XENT baseline in all cases, specifically for a given model training method we report the difference between its performance and the XENT baseline performance. This does not change that a lower value is better for the APCER metric.
Our final ranking values are averages over all attack types compiled from multiple random variables: the variable for multiple runs per attack type left out, and the average over each category. Due to this we simply report point value ranking and exclude standard deviations. We define a better performance as lower (in the case of APCER).
@@ 3.2 Dataset Sources
The images in our dataset are the same set used in Boyd et al. [1] and Dowling et al. [15] experiments and reflect the following attack types and anomalies along with the bona fide samples and their sources: 198 real irises [6, 34, 18, 60, 49, 33, 69, 66, 68], 83 irises with textured contact lenses [34, 33, 69, 68], 57 artificial (plastic eyes and glass prosthesis) eyes [34], 95 diseased/unhealthy eyes [60, 59], 91 printouts [34, 18, 32], 89 synthetically-generated iris images [67], 87 printouts of irises with textured contact lenses [34], and 65 post-mortem iris images [63].
The eye tracking heatmaps were sourced from an eye tracking study conducted by Dowling et al. [15] who captured eye gaze heatmaps from a mixture of expert and non-expert examiners. Acquiring this variety of saliency allows us to perform the first study we are aware of comparing the performance of different human saliency types for saliency-guided training in the task of iris PAD.
@@
3.3 Saliency Methods
3.3.1 Segmentation Masks
3.3.2 Hand Annotations
3.3.3 Eye Tracking
Raw eye tracking data collected from both expert and non-expert participants during iris image evaluations was processed into individual saliency heatmaps. For each participant normalized fixation coordinates were first corrected using per-participant polynomial remapping coefficients to account for calibration offset, and a Gaussian Blur was applied to produce a continuous heatmap representation of visual attention. To produce a unified ground truth saliency map for each iris image, individual participant heatmaps were aggregated via a pixel-wise mean across all evaluators, expert and non-expert.
3.4 No Saliency: Foundation Models
Research has shown that foundation models show a remarkable ability to distinguish both full-size and cropped iris samples in closed-set presentation attack detection settings [53, 54]. Specifically, Redwan et al. used a variety of foundation models to obtain feature embeddings for iris PAD images and trained three standard classifiers, logistic regression (LogReg) [8], support vector machine (SVM)[7], and support vector machine with radial basis function kernel (SVM-RBF) [50] on the extracted features. In the task of iris PAD for full size irises, [42] suggests that the clear winner was DINOv2. Motivated by these findings, we selected DINOv2-Large and the newer variant DINOv3 [52] to the list of models to test. Finally, motivated by the findings of Piland et al. [46], whose experiments found that SigLIP2 [64] pre-trained embeddings showed good clustering of iris PAD attack types, we select this as our final foundation model to test. We test open-set iris PAD using similar methodology laid out by Redwan et al. by training three standard classifiers, SVM-Linear, SVM-RBF, and a Spatial-CNN, with the embeddings from these three foundation models. Although fine-tuning models can lead to incremental gains in peak performance for specific tasks, we intentionally employ these models as frozen feature extractors. This approach is chosen to ensure consistency with the methodologies established in prior work, avoid high computational overhead, and prevent ’feature drift’ from overwriting the general capabilities essential for open-set robustness [58, 72].
Table 1: Compact $\Delta$ APCER @ BPCER=1% vs XENT Baseline ($\downarrow$) results are presented.
@@ 3.4 No Saliency: Foundation Models
High Entropy +0.0256 +0.0088 +0.0167 +0.0169 +0.0623 -0.0293 +0.0000 +0.0144
Equal Entropy +0.0879 -0.0535 +0.0423 +0.0094 +0.0172 -0.0340 +0.0088 +0.0112
Low Entropy -0.0403 +0.0079 +0.0282 +0.0094 +0.1676 -0.0082 +0.0102 +0.0250
@@
Full Eye Tracking -0.2766 -0.0211 +0.0154 -0.0665 +0.0824 -0.1197 -0.2354 -0.0888
Initial Eye Tracking -0.2701 -0.0474 +0.0744 -0.0431 +0.1753 -0.0833 -0.3816 -0.0823
De-noised Full ET -0.2894 -0.0412 +0.0718 -0.0927 +0.2193 -0.1127 -0.3553 -0.0857
De-noised Initial ET -0.3617 -0.0412 +0.0551 -0.1414 +0.1600 -0.0599 -0.3553 -0.1063
@@
DINOv2-Large + CNN +0.3724 +0.0037 +0.1010 -0.0576 +0.6388 -0.0073 -0.0127 +0.1483
DINOv2-Large + SVM-Linear +0.4027 -0.0056 +0.0892 -0.0373 +0.6582 +0.0043 +0.0018 +0.1590
DINOv2-Large + SVM-RBF +0.4191 +0.0268 +0.0940 -0.0685 +0.6966 +0.0479 -0.0157 +0.1715
DINOv3 + CNN +0.3031 -0.1312 -0.1780 -0.1594 +0.3539 -0.2010 -0.0617 -0.0106
DINOv3 + SVM-Linear +0.2428 -0.1047 -0.1496 -0.2814 +0.3251 -0.0549 -0.0169 -0.0057
DINOv3 + SVM-RBF +0.4071 -0.0649 -0.0927 -0.2330 +0.5127 -0.0156 -0.0084 +0.0721
SigLIP2 + CNN +0.3216 -0.3820 -0.1160 -0.1779 +0.4142 -0.0237 -0.0666 -0.0043
SigLIP2 + SVM-Linear +0.2398 -0.2085 -0.1428 -0.1841 +0.6245 +0.0042 -0.0422 +0.0416
SigLip + SVM-RBF +0.3171 -0.1531 -0.2068 -0.1452 +0.6574 -0.2049 -0.0467 +0.0311
@@
Table 2: Comparison of Saliency-Augmented Training and Foundation Model Performance.
(a) Pooled APCER at 1% Threshold (Random-Effects)
Rank Method Pooled 95% CI $I^{2}$
1 De-noised Initial ET 0.6765 [0.4182, 0.9349] 97.8%
2 Full Eye Tracking 0.6918 [0.4385, 0.9450] 97.2%
3 De-noised Full ET 0.6968 [0.4629, 0.9306] 98.2%
4 Initial Eye Tracking 0.7007 [0.4564, 0.9450] 98.4%
5 DINOv3 + CNN 0.7680 [0.6576, 0.8784] 98.8%
6 DINOv3 + SVM-Linear 0.7731 [0.6334, 0.9127] 99.3%
7 SigLIP2 + CNN 0.7756 [0.6484, 0.9029] 98.5%
8 Densenet + Baseline 0.7946 [0.5416, 1.0477] 98.1%
9 SigLIP2 + SVM-RBF 0.8107 [0.7100, 0.9115] 98.9%
10 SigLIP2 + SVM-Linear 0.8214 [0.7341, 0.9088] 98.8%
11 Hand Annotations Equal Entropy 0.8344 [0.6191, 1.0497] 97.9%
12 DINOv3 + SVM-RBF 0.8571 [0.7611, 0.9531] 99.3%
13 Segmentation Masks 0.8677 [0.6774, 1.0580] 97.6%
14 Hand Annotations Low Entropy 0.8769 [0.7033, 1.0505] 97.5%
15 Hand Annotations High Entropy 0.8936 [0.7316, 1.0556] 96.9%
16 DINOv2-Large + CNN 0.9264 [0.8955, 0.9572] 99.3%
17 DINOv2-Large + SVM-Linear 0.9372 [0.9009, 0.9735] 99.0%
18 DINOv2-Large + SVM-RBF 0.9493 [0.9266, 0.9720] 99.6%
(b) Robustness Summary (Wins/Losses)
Method Sig+ Sig– Avg $|d|$
DINOv3 + CNN 5 2 2.469
SigLIP2 + SVM-RBF 5 2 3.099
De-noised Initial ET 4 0 1.180
DINOv3 + SVM-Linear 4 2 2.105
SigLIP2 + CNN 4 2 2.551
SigLIP2 + SVM-Linear 4 2 2.255
Initial Eye Tracking 3 0 1.276
De-noised Full ET 3 0 1.188
Full Eye Tracking 3 0 0.854
DINOv3 + SVM-RBF 1 2 2.081
DINOv2-Large + SVM-Linear 1 2 1.503
DINOv2-Large + CNN 1 3 1.801
DINOv2-Large + SVM-RBF 1 3 2.191
Hand Annotations Equal Entropy 0 0 0.395
Hand Annotations Low Entropy 0 0 0.318
Segmentation Masks 0 0 0.437
Hand Annotations High Entropy 0 1 0.265
(Methods ranked by significant improvement)
@@ 4 Results
4.1 Statistical Analysis
For each configuration, 12 independent models per attack-type-left-out were trained for a total of 84 models tested. For each run of 12 models, we report the average APCER at BPCER 1% threshold. To properly contextualize our results, we first checked Pearson’s Correlation Coefficient for all methods against the XENT baseline and confirmed the correlation was near zero for all methods indicating statistical independence. We then employed Welch’s unpaired t-test ($p<0.05$) for all subsequent comparisons to account for unequal variance. Next, we conducted per-attack comparisons of each model configuration against the XENT baseline and using Cohen’s d effect size, we identify both significant improvement and significant degradation. To evaluate how the results generalize across the entire attack battery, we performed a random-effects meta-analysis using DerSimonian-Laird estimator for between-attack heterogeneity ($\tau^{2}$). To ensure conservative and reliable estimates given the number of attack categories ($k=7$), we applied the Hartung-Knapp-Sidik-Jonkman (HKSJ) adjustment to the 95% confidence interval. Methods in table 2 are ranked by their pooled random-effects accounting for within-attack variance and between-attack heterogeneity.
@@ 4.1 Statistical Analysis
Table 1 shows AP/BPCER 1% results expressed as each method’s improvement above the XENT baseline where each row refers to a method and each column refers to the attack type that was left out during training with a final delta column summarizing the overall performance. Our findings indicate that the De-noised Initial ET configuration shows the largest improvement over the XENT baseline with an improvement of 0.1063. The remaining three eye tracking saliency configurations, Full Eye Tracking, De-noised Full ET, and Initial Eye Tracking, improved the XENT baseline by 0.0888, 0.0857, and 0.0823 respectively. The remaining saliency-based methods, hand annotations and segmentation masks, failed to surpass the XENT baseline entirely. These results are further reinforced when performing meta-analysis in table 2. Ranking by weighted mean to account for variance, pooled AP/BPCER 1% results indicate that eye tracking saliency methods outperformed all other saliency methods with De-noised Initial ET performing the best of all configurations with a pooled average of 0.6765 with the other eye tracking methods closely behind. Thus the answer to RQ1 is affirmative, eye tracking saliency, particularly De-noised Initial Eye Tracking saliency, surpasses other saliency-based methods in open-set iris PAD.
@@
When comparing the results of foundation models in 1, only three of the nine model configurations were able to surpass the XENT baseline, DINOv3 + CNN, DINOv3 + SVM-Linear, and SigLIP2 + CNN. When applying meta-analysis to these results in 2, the weighted average improves with pooled averages of 0.7680 for DINOv3 + CNN, 0.7731 for DINOv3 + SVM-Linear, and 0.7756 for SigLIP2 + CNN respectively compared to a pooled average of 0.7946 for XENT. For saliency-based methods, segmentation masks and all hand annotation configurations failed to surpass the pooled average AP/BPCER 1% of XENT. Eye tracking methods however show strong improvement over XENT with the top four performing configurations for pooled AP/BPCER belonging to an eye tracking method and De-noised Initial ET recording the best pooled AP/BPCER at 0.6765. With eye tracking methods improving XENT APCER by 0.0978 or greater and only three foundation model configurations improving the baseline at all, the answer to RQ2 is clear, saliency-guided training using eye tracking heatmaps outperform foundation models in open-set iris PAD.